# TractoGPT: A GPT architecture for White Matter Tract Segmentation

### v1

@@ 1 Introduction
Fiber tract segmentation is a pivotal process in Neuroimaging, enabling detailed analysis of White Matter connectivity through Diffusion Magnetic Resonance Imaging (dMRI). Tractography traces the anisotropic diffusion of water molecules along neural pathways, yielding three-dimensional streamlines that represent white matter fiber bundles. These streamlines are grouped into specific anatomical tracts, providing crucial insights into brain connectivity and function, essential for understanding development, aging, and neurological conditions [1]. With recent advancements, fiber tract segmentation methods can be broadly categorized into classical and deep learning techniques. Classical methods, such as QuickBundles and QuickBundlesX, use Mean Direct-Flip Distance (MDF) to cluster streamlines into bundles or tracts [2]. Other Techniques, like WMA [3], incorporate anatomical priors, while RecoBundles [4] recognizes model bundles, and BundleSeg [5] enhances this recognition within target tracts via Fast Streamline Search. Deep learning Tract Segmentation approaches, for instance, TractSeg [6] employs Convolutional Neural Networks (CNNs) across multiple MRI slices, and DeepWMA [7] utilizes novel fiber descriptors. Additionally, clustering techniques like CINTA [8] leverage streamline clustering, and FINTA [9] does filtering in embedding space, and FIESTA [10] further improves the performance by employing autoencoders to segment and filter streamlines in latent space through FINTA multibundles, then generating additional streamlines with GESTA-GMM [11] to fill bundles and meet semi-automatically calibrated bundle-specific thresholds. Recent tract segmentation studies have also explored point cloud networks [12, 13, 14]. But many methods stated above either require registration, ATLAS, filtering or calibration for thresholds. In this paper,
  • We introduce TractoGPT, a novel, fully-automatic, registration-free white matter segmentation network inspired by the GPT architecture using auto-regressive pretraining.
  • In addition to TractoGPT framework, we introduce a
@@ 2 Methodology
Fig. 1: TractoGPT Architecture: (Stage I) Raw Streamline undergoes preprocessing (i, ii, iii, iv) to give us 3 different data representations, refer Section 2.2. (Stage II) From 3 different point cloud arrays, any one can be chosen to train TractoGPT, Extracted Point Cloud undergoes FPS (Farthest Point Sampling) to give total $P$ center points, which are used to sample a total of $K$ nearest neighbors using kNN (k-nearest neighbors) (v), creating point cloud patches. (Stage III) Point patches get sequence using Morton order (vi), and a PointNet-style encoder gives embedding for each patch which are fed as tokens to the Transformer model.
@@ 2.2 Data Representations
To embed a richer understanding of tractography data to the model, we propose 3 data representations, streamline, cluster, and fusion data (refer Fig. 1, Stage I), either of which can be used as the Training Data. Multiple data representations are used to combat local information deficiency in single streamline data.
@@
    (Note: Sampled Streamlines in this method are not interpolated and clusters are mutually exclusive)
  • Fusion: Fusion Data is a fusion of streamline and cluster data allowing the amalgamation of both representations. For fusion data, 256 points are sampled from the interpolated streamline of interest, and the rest 768 points are sampled from the non-interpolated neighboring streamlines to make a 1024-dimensional array (see (iii), (iv) in Fig.1).(Note: Interpolation allows focus on selected streamline, and diffused focus on neighbours
Due to time efficiency, and bounded parent clusters, we use QuickBundlesX for finding neighbouring streamlines. Based on the choice of data representation, see Section 2.2, our Input Point Cloud data, X, is either of shape $(256,3)$ in streamline representation or $(1024,3)$ in Fusion and Cluster Representations. As Point Cloud data transitions from Stage I to Stage II, refer Fig. 1, also point cloud arrays are shuffled.
@@ 2.3 Tokenization
Points in point clouds individually are spatial coordinates and have minimal information to learn spatial representations, whereas a group of points give regional information on the shape of the point cloud. We create point patches to embed regional information in tokens using an encoder network.
As mentioned in the caption of Fig. 1, point patches are obtained through FPS-kNN (Farthest Point Sampling & k-Nearest Neighbors) where the farthest points are treated as centroids to make patches by sampling K neighbors via kNN. For GPT architecture, sequential information is required among tokens which are extracted using sorted Morton Order (or Z-order curve) on encoded $K$ center points (1-dimensional array), and this Relative Positional Encoding is passed to the Generator part of TractoGPT (see Fig. 1)[17]. The coordinates of each point are normalized relative to its center point before they are fed to the PointNet-style encoder [18], giving a latent representation per patch to make tokens, along with the Morton order sequence. While setting patch configuration, a patch size of 32 points for [Fusion, Cluster] and 8 points for Streamline. Number of patches is set to 64 for all representations to ensure overlapping patches, maximizing retention of continuous information across the whole input point cloud X.
@@ 2.4 TractoGPT Model
In TractoGPT, sequence of tokens extracted from the Tokenization process (refer to Section 2.3) undergo autoregressive pre-training strategy on all data representations separately (refer Fig. 1), ablation across these representations have been done in Table 2. We propose an architecture consisting Extractor and Generator which are essentially stacked transformer decoder blocks, inspired by PointGPT, Point-BERT, and other point cloud architectures [17, 19, 20].
Overall pre-training objective of TractoGPT is to reconstruct input point cloud patch X coordinates learning latent representations which would be later fine-tuned for downstream tasks. While pre-training we use dual masking strategy which does intermittent masking on top of causal masking inhibiting model to overfit on the input point cloud data. Causal masking is used in training Next Token Prediction tasks and intermittent masking can be understood as a proportion of masked preceding tokens attending to each unmasked preceding token.
Due to random transformations of points while training, order of points is not preserved, leading to ambiguity in predicting consecutive patches. To mitigate this ambiguity the Generator incorporates directions as Relative Patch Positions Encoding (refer Fig. 1 Stage II and III), without disclosing the locations of masked patches or the overall shapes of the point clouds. The extractor takes sinusoidal positional encoded tokens sorted by Morton ordering, prediction head of the Generator is designed to predict subsequent point patches in the coordinate space. The Generator comprises of 2 FCNN layers with (ReLU) activation functions, and is shallower than the Extractor.
Losses: Goal of the Generator is to reconstruct patch coordinates from patch tokens. For pre-training, generation loss is calculated using summation of L1-norm (CDL1) and L2-norm (CDL2) of the Chamfer Distance (CD) in equal proportions. During fine-tuning, we freeze the Generator and fine-tune the Extractor on cross-entropy loss along with (CDL1, CDL2) for the downstream classification task to prevent overfitting and increase robustness.
@@ 2.5 Model Training and Testing
A consistent number of streamlines per class across all training subjects is used to mitigate imbalance for training TractoGPT (for example: we use 500 streamlines per subject per class on a single streamline which can vary based on the choice of data representation). But while testing, all streamlines of a test subject are classified, without leaving a single streamline behind. In TractoGPT training and testing strategy, we implement a comprehensive approach that includes pretraining, fine-tuning, and testing. Pretraining involves reconstruction of patch coordinates using a 50:50 Chamfer Loss (L1 and L2 norm) without labels, optimized with AdamW (weight decay of 0.05) and a cosine learning rate scheduler starting at 0.0001 learning rate, over a maximum of 100 epochs (converges earlier). In fine-tuning for classification, we employ a combination of Cross Entropy and Chamfer Distance Loss (CDL1 + CDL2). The overall strategy is designed to optimize the model’s performance in classification and reconstruction tasks while leveraging advanced loss functions for effective tractography streamlines understanding. TractoGPT takes about 4 days on average to converge (cluster $<$ streamline $<$ fusion) and 12 hours to infer on a single V100 16 GB respective splits of TractoInferno Dataset.
@@ 3 Experiments and Results
We perform rigorous experiments with the current state-of-the-art tractography segmentation models, FINTA-m, RBx across the list of 23 common tracts in TractoInferno and HCP dataset. We show the Ablation Study, in the Table 2, where we can see Fusion is comparable to the Cluster, proving the efficacy of Fusion Representation at large. For Comparative Study, we choose TractoGPT on cluster data representation as it performs better than other methods. We compare TractoGPT [cluster] with SOTA methods as shown in Figure 2 where we can see our method segments all common 23 major bundles with good DICE scores. Here, RecoBundlesX outputs are not filtered using dMRIQC [21]. We also demonstrate TractoGPT’s generalization capability, by training TractoGPT on HCP dataset with 72 classes, i.e. TractoGPT-hcp, and testing on TractoInferno test data for which results are shown in Table 3. TractoGPT-hcp indicates better generalization than FIESTA. FIESTA reports DICE $0.74\pm 0.08$ on the private Myeloinferno dataset [10], while TractoGPT-hcp achieved $0.79\pm 0.12$ on TractoInferno. We only provide results on DICE scores to provide a fair comparison of a common metric; other metrics like Hausdorff distance are not used in these works.

### v2

@@ 1 Introduction
Fiber tract segmentation is a pivotal process in Neuroimaging, enabling detailed analysis of White Matter connectivity through Diffusion Magnetic Resonance Imaging (dMRI). Tractography traces the anisotropic diffusion of water molecules along neural pathways, yielding three-dimensional streamlines that represent white matter fiber bundles. These streamlines are grouped into specific anatomical tracts, providing crucial insights into brain connectivity and function, essential for understanding development, ageing, and neurological conditions [1].
With recent advancements, fiber tract segmentation methods can be broadly categorized into classical and deep learning techniques. Classical methods, such as QuickBundles and QuickBundlesX[2] cluster streamlines into bundles, Fast Streamline Search [3] searches similar streamlines accurately, RecoBundles[4] recognizes model bundles, utilizing distance metrics like Mean Direct-Flip Distance (MDF)[2]. Deep Learning techniques like TractSeg[5] employs Convolutional Neural Networks (CNNs) across multiple MRI slices, DeepWMA [6] utilizes novel fiber descriptors, FINTA [7] does filtering in embedding space, and FIESTA [8] further improves the FINTA by employing FINTA-multibundle to segment and GESTA-GMM [9] to fill bundles to meet semi-automatically calibrated bundle-specific thresholds. Recent tract segmentation studies have also explored point cloud networks [10, 11], but most methods either require registration, ATLAS, filtering or calibration for thresholds. In this paper,
  • We introduce TractoGPT, a novel, fully-automatic, registration-free white matter segmentation network inspired by the GPT architecture.
  • In addition, we introduce a
@@ 2 Methodology
Fig. 1: TractoGPT Architecture: (Stage I) Raw Streamline undergoes preprocessing (i, ii, iii, iv) to give us 3 different data representations (Section 2.2). (Stage II) From 3 different point cloud arrays, any one can be chosen to train TractoGPT, Extracted Point Cloud undergoes FPS (Farthest Point Sampling) to give total $P$ center points (Absolute Positions), used to sample a total of $K$ nearest neighbors using kNN (v), creating point cloud patches. (Stage III) Point patches get sequence using Morton order (Relative Positions) (vi), and a PointNet-style encoder gives embedding for each patch as tokens.
@@ 2.2 Data Representations
To embed a richer understanding of tractography data to the model, we propose 3 data representations, (refer Fig. 1, Stage I), either of which can be used as the Training Data. Due to time efficiency, and bounded parent clusters, we use QuickBundlesX for finding neighbouring streamlines.
@@
  • Fusion: Fusion Data is a fusion of streamline and cluster data allowing the amalgamation of both representations. For fusion data, 256 points are sampled from the interpolated streamline of interest, and the rest 768 points are sampled from the non-interpolated neighboring streamlines to make a 1024-dimensional array (see (iii), (iv) in Fig.1).
@@ 2.3 Tokenization
A group of points yields regional information on the shape of the point cloud. We create point patches to embed regional information in tokens using an encoder network.
Point patches are obtained through FPS-kNN (Farthest Point Sampling & k-Nearest Neighbors) where the farthest points are treated as centroids to make patches by sampling K neighbors via kNN. For GPT architecture, sequential information is required among tokens which are extracted using sorted Morton Order (or Z-order curve) on encoded $K$ center points (1-dimensional array), and this Relative Positional Encoding is passed to the Generator part of TractoGPT (see Fig. 1)[14]. The coordinates of each point are normalized relative to its center point before they are fed to the PointNet-style encoder [15], giving a latent representation per patch to make tokens, along with the Morton order sequence. While setting patch configuration, a patch size of 32 points for [Fusion, Cluster] and 8 points for Streamline. Number of patches is set to 64 for all representations to ensure overlapping patches.
@@ 2.4 TractoGPT Model
In TractoGPT, we employ an architecture consisting of Extractor and Generator which are essentially stacked transformer decoder blocks [14, 16]. The model undergoes autoregressive pretraining separately on all data representations (refer Fig. 1). The overall pre-training objective of our model is to reconstruct input point cloud patches using sequence of tokens extracted from the Tokenization process above. In pre-training we use dual masking strategy which does intermittent masking on top of causal masking inhibiting model to overfit on the input point cloud data. Causal masking is usually used for Next Token Prediction tasks, and intermittent masking can be understood as a proportion of masked preceding tokens attending to each unmasked preceding token.
Due to random transformations and shuffling of points while training, order of points is not preserved, leading to ambiguity in predicting consecutive patches. To mitigate this ambiguity, Extractor takes Sinusoidal Positional encoded tokens with absolute normalized positions of center points, and the Generator incorporates directions as Relative Patch Positions Encoding (refer Fig. 1 Stage II and III), without disclosing the locations of masked patches or the overall shapes of the point clouds. The Prediction head of the Generator is designed to predict subsequent point patches in the coordinate space. The Generator comprises of 2 FCNN layers with (ReLU) activation functions, being shallower than the Extractor.
@@ 2.5 Model Training and Testing
A consistent number of streamlines per class across all training subjects is used to mitigate imbalance for training TractoGPT (for example: we use 500 streamlines per subject per class on a single streamline data representation which can vary based on the choice of data representation). But while testing, all streamlines of a test subject are classified, without leaving a single streamline behind. In TractoGPT training and testing strategy, we implement a comprehensive approach that includes pretraining, fine-tuning, and testing. Pretraining involves reconstruction of patch coordinates using a 50:50 Chamfer Loss (L1 and L2 norm) without labels, optimized with AdamW (weight decay of 0.05) and a cosine learning rate scheduler starting at 0.0001 learning rate, over a maximum of 150 epochs (converges earlier). In fine-tuning for classification, we employ Cross Entropy and Chamfer Distance Loss (CDL1 + CDL2) in a ratio of 1:3 respectively. The overall strategy is designed to optimize the model’s performance in classification and reconstruction tasks while leveraging advanced loss functions for effective tractography streamlines understanding. TractoGPT takes less than 4 days on average to converge (cluster $<$ streamline $<$ fusion) and 12 hours to infer on a single V100 16 GB GPU on TractoInferno Dataset.
@@ 3 Experiments and Results
We perform rigorous experiments with the current state-of-the-art tractography segmentation models, FINTA-m, RBx across the list of 23 common tracts in TractoInferno and HCP dataset. For a fair and consistent comparison across methods we use DICE scores. We show the Ablation Study, in the Table 2, where we can see Fusion is comparable to the Cluster, proving the efficacy of Fusion Representation at large. For the comparative study (see 3), we chose TractoGPT in cluster data representation because it performs better than other methods on TractoInfero dataset. We show tract-wise comparison of TractoGPT [cluster] with other SOTA methods as shown in Figure 2 where we can see that our method segments all 23 common major bundles with good DICE scores. Here, RecoBundlesX outputs are not filtered using dMRIQC [17]. We also demonstrate TractoGPT’s generalization capability, by training TractoGPT on HCP dataset and testing on TractoInferno test data (see resultes in Table2), we named it TractoGPT-hcp. TractoGPT-hcp indicates better generalization than FIESTA. FIESTA reports DICE score of $0.74\pm 0.08$ on a private dataset, Myeloinferno[8], whereas TractoGPT-hcp achieved $0.79\pm 0.12$ on TractoInferno.